OpenStudy (anonymous):

The new product is a pyramid-shaped gum with a square base. In the spirit of the other challenges, the company has decided to place their pyramid-shaped gum inside a clear, glass, giant, bubblegum-shaped sphere. Each piece of gum has a base measurement of 1 inch and a height of 0.75 inches. The glass sphere container has a diameter of 17.25 inches. How many pieces of Pharaoh Chewing Gum can fit inside the glass sphere?

OpenStudy (anonymous):

@sourwing @surjithayer

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@mathstudent55 help please!

OpenStudy (mathstudent55):

I have no idea. Sorry.

OpenStudy (anonymous):

@mathstudent55 ok thanks though, np

OpenStudy (anonymous):

@surjithayer anything? :(

OpenStudy (anonymous):

volume of one pyramid\[=\frac{ 1 }{ 3 }*base^2*height=\frac{ 1 }{ 3 }*1^2*0.75=0.25 ~inch^3\] radius of sphere\[=\frac{ diameter }{ 2 }=\frac{ 17.25 }{ 2 }=8.625~inch\] volume pf sphere\[=\pi *r^3=3.14*\left( 8.625 \right)^3 ~inch^3\] \[no.~of~gums~contained~one~sphere=\frac{ 3.14*\left( 8.625 \right)^3 }{ 0.25 }=?\]

OpenStudy (anonymous):

Ok one sec, let me calculate

OpenStudy (anonymous):

40293.68203

OpenStudy (anonymous):

@surjithayer

OpenStudy (mathstudent55):

If the gum were a liquid, then you can calculate the volume of one piece of gum, and calculate the volume of the sphere, and divide the volume of the sphere by the volume of one piece of gum to calculate how many "liquid pieces of gum" fit in the sphere, but the gum is a solid with a certain shape. As pieces of gum fall into the sphere, they will leave unoccupied air space. Not every cubic inch of volume of the sphere will be filled. That's why I don't know how to calculate this.

OpenStudy (anonymous):

oh alright, thanks you and no problem :)) @mathstudent55